# Fast LiDAR Data Generation with Rectified Flows

### v1

@@ IV-A Settings
Dataset. Following baselines [5, 6, 4], we utilize the KITTI-360 [16] dataset. The KITTI-360 dataset contains 81,106 point clouds captured using a Velodyne HDL-64E (64-beam mechanical LiDAR sensor). We adopt the standard data split defined by Zyrianov et al. [5]. Each point cloud is projected onto a $64\times 1024$ image, with range and reflectance values assigned to each pixel.
@@
LiDM w/o APE LiDM w/ APE Dataset
Figure 4: Distribution of point clouds generated by LiDM [6]. We calculated a marginal distribution of 1000 random samples. With APE, the distribution get closer to the dataset.
Evaluation metrics. Following the related work, we evaluate the distributional similarity between real and generated samples across multiple levels of data representation. We use seven evaluation metrics: Fréchet range distance (FRD)[5], Fréchet range image distance (FRID)[6], Fréchet point cloud distance (FPD)[30], Fréchet point-based volume distance (FPVD)[6], Fréchet sparse volume distance (FSVD)[6], Jensen–Shannon divergence (JSD)[5], and maximum mean discrepancy (MMD) [5], based on five types of data representations: range images, reflectance images, point clouds, voxels, and bird’s-eye views (BEV). The point cloud, voxel, and bird’s-eye view representations are derived from the range image. Table I summarizes the correspondence between the metrics and the representations. Note that only FRD incorporates both range and reflectance modalities, while the other metrics rely solely on the range modality. For each method, we generate 10,000 samples and evaluate them against the entire dataset of real samples, following standard practices in generative models.
Implementation details. All models are implemented using PyTorch. We performed a distributed training with automatic mixed precision (AMP). We used torchdiffeq [31] for solving ODEs. For training 2-RF, we sampled 1M pairs by the dopri5 sampler (adaptive step-size) with absolute/relative tolerance of 1e-5. For timestep distillation, we sampled 100k pairs with the same sampler. All evaluation results are produced by the euler sampler (fixed step-size) for fair comparison. Our code and pretrained weights will be available on https://github.com/kazuto1011/r2flow.
@@ IV-B Results
Quantitative results. Tab. I shows the evaluation results in three groups from top to bottom In the first group which compares GAN-based methods, DUSty v2 that generates both range and reflectance images performed well across multiple metrics. The second group compares the results of the iterative models including ours, with higher NFEs (number of function evaluations). NFE counts the number of times running neural networks. With the introduction of APE, LiDM demonstrates improvements across all metrics, particularly in BEV-based JSD and MMD. Our R2Flow achieved results comparable to another pixel-space model, R2DM. On the other hand, the 2-RF scores were slightly lower, which is because the quality upper bound for 2-RF is limited by the parent model 1-RF rather than real data. Overall, the iterative models are better than the GANs. The third group compares the results with fewer NFEs. While all baseline methods exhibit significant performance degradation, our R2Flow mitigates the degradation through reflow and distillation. By incorporating the few-step distillation (2-TD and 4-TD), some metrics exhibit results comparable to those achieved with a larger number of steps. Fig. 5 shows FRD scores as a function of NFEs in details. R2Flow shows the better computational tradeoff.
@@
  • Notation: range image, reflectance image, point cloud, voxel, bird’s eye view.
@@
Figure 5: Speed–quality tradeoff in FRD. We compare the methods [2, 3, 5, 4] supporting both range and reflectance modalities. Our R2Flow (blue lines) shows the better tradeoff against the baselines (black lines).
Qualitative results. In Fig. 6, we compare generated LiDAR point clouds from real data [16], DUSty v2 [3], LiDARGen [5], our improved LiDM [6], R2DM [4], and our R2Flow. LiDM, R2DM, and R2Flow demonstrate better quality, such as sharper scan lines and clearer object boundaries, while LiDM exhibits some wavy and blurry boundaries, as previously reported by Ran et al. [6].
@@
Figure 7: Trajectory curvature of learned flows. (a) Trajectory curvatures between 1-RF and 2-RF (b) top-200 curved trajectories (0.15% of all pixels). The pixel value $-1$ at $t=1$ corresponds to raydrop noise.
@@ V Conclusions
In this paper, we presented R2Flow, the rectified flow-based generative model for fast and realistic LiDAR data generation. We verified the effectiveness of our approach in both efficiency and quality in the unconditional generation evaluation. Future work will focus on investigating the scalability of R2Flow, improving the reflow process to maintain quality, and demonstrating R2Flow in application tasks such as sparse-to-dense completion, sim-to-real domain adaptation, and anomaly detection. The trajectory visualization in Fig. 7 suggests that the raydrop pixels drifting toward a value of $-1$ may hinder training straight flows. We expect that implementing a raydrop-aware architecture [2, 3] could mitigate this issue.

### v2

@@ IV-A Settings
Dataset. Following prior work [5, 6, 4], we utilize the KITTI-360 [17] dataset. The KITTI-360 dataset contains 81,106 point clouds captured using a Velodyne HDL-64E (64-beam mechanical LiDAR sensor). We adopt the standard data split defined by Zyrianov et al. [5]. Each point cloud is projected onto a $64\times 1024$ image with range and reflectance values assigned to each pixel.
@@
LiDM LiDM w/ APE Dataset
Figure 5: Distribution of point clouds in bird’s eye view. We calculated the marginal distribution of 1,000 random samples generated by LiDM [6]. With APE, the distribution gets closer to the dataset.
Evaluation metrics. Following the related work, we evaluate the distributional similarity between real and generated samples across multiple levels of data representation. We use seven evaluation metrics: Fréchet range distance (FRD)[5], Fréchet range image distance (FRID)[6], Fréchet point cloud distance (FPD)[32], Fréchet point-based volume distance (FPVD)[6], Fréchet sparse volume distance (FSVD)[6], Jensen–Shannon divergence (JSD)[5], and maximum mean discrepancy (MMD) [5], based on five types of data representations: range images, reflectance images, point clouds, voxels, and bird’s-eye views (BEV). The point cloud, voxel, and BEV representations are derived from the range image. Table I summarizes the correspondence between the metrics and the representations. Note that only FRD incorporates both range and reflectance modalities, while the other metrics rely solely on the range modality. For each method, we generate 10,000 samples and evaluate them against the entire dataset of real samples, following standard practices in generative models.
Implementation details. All models are implemented using PyTorch. Training and evaluations were performed on four NVIDIA RTX 6000 Ada GPUs. We performed a distributed training with automatic mixed precision (AMP). We used torchdiffeq [33] for solving ODEs. For training 2-RF, we sampled 1M pairs by the dopri5 sampler (adaptive step-size) with absolute/relative tolerance of 1e-5. For timestep distillation, we sampled 100k pairs with the same sampler. All evaluation results are produced by the euler sampler (fixed step-size) for fair comparison. Our code and pretrained weights are available at https://github.com/kazuto1011/r2flow.
@@ IV-B Results
Quantitative results. Tab. I shows the evaluation results in three groups from top to bottom In the first group which compares GAN-based methods, DUSty v2 that generates both range and reflectance images performed well across multiple metrics. The second group compares the results of the iterative models including ours, with a higher NFE (number of function evaluations). NFE counts the number of times running neural networks. With the introduction of APE, LiDM demonstrates improvements across all metrics, particularly in BEV-based metrics, JSD and MMD. Our R2Flow achieved results comparable to another pixel-space model, R2DM. On the other hand, the 2-RF scores were slightly lower, which is because the quality upper bound for 2-RF is limited by the parent model 1-RF rather than real data. Overall, the iterative models are better than the GANs. The third group compares the results with a fewer NFE. While all baseline methods exhibit significant performance degradation, our R2Flow mitigates the degradation through reflow and distillation. By incorporating the few-step distillation (2-TD and 4-TD), some metrics exhibit results comparable to those achieved with a larger number of steps. Fig. 6 shows FRD scores as a function of NFEs in details. R2Flow shows a better computational tradeoff.
@@
  • Notation: range image, reflectance image, point cloud, voxel, bird’s eye view (BEV).
@@
Figure 6: Speed–quality tradeoff in FRD. We compare the methods [2, 3, 5, 4] that support both range and reflectance modalities. Our R2Flow (blue lines) shows the better tradeoff against the baselines (black lines).
Qualitative results. In Fig. 7, we compare LiDAR point clouds generated from DUSty v2 [3], LiDARGen [5], our improved LiDM [6], R2DM [4], and our R2Flow. LiDM, R2DM, and R2Flow demonstrate better quality, such as sharper scan lines and clearer object boundaries, while LiDM exhibits some wavy and blurry boundaries, as previously reported by Ran et al. [6].
@@
Figure 8: Trajectory curvature of learned flows. (a) Trajectory curvature between 1-RF and 2-RF (b) top-200 curved trajectories (0.15% of all pixels). The pixel value $-1$ at $t=1$ corresponds to raydrop noise.
@@ V Conclusions
In this paper, we presented R2Flow, the rectified flow-based generative model for fast and realistic LiDAR data generation. We verified the effectiveness of our approach in both efficiency and quality in the unconditional generation evaluation. Future work will focus on exploring the scalability of R2Flow, refining the reflow process to maintain quality, and demonstrating its effectiveness in application tasks such as sparse-to-dense completion, sim-to-real domain adaptation, and anomaly detection. The trajectory visualization in Fig. 8 suggests that the raydrop pixels drifting toward a value of $-1$ may hinder the training of straight flows. We anticipate that implementing a raydrop-aware architecture [2, 3] could mitigate this issue.